OpenStudy (anonymous):

3(2x+1)-4(x-2)=10 Please can you show it step by step as well.

OpenStudy (anonymous):

\[6x+3-4x+8=10\] \[2x+11=10\] \[2x=-1\] \[x=-1/2\]

OpenStudy (across):

\[3(2x+1)-4(x-2)=10\]\[6x+3-4x+8=10\]\[2x+11=10\]\[2x=-1\]\[x=-1/2\]mathematicalpro is correct.